if 15% is 200.00 then how much is 85%?
Found 2 solutions by colliefan, Alan3354:
Answer by colliefan(242)

Percentage Discussion
For all percentage problems, remember that the basic equation or formula is that the percentage is a ratio of some “part” divided by some “whole” or standard amount. That doesn't seem precise, but I think that language is helpful for all percentage problems. By part and whole, I mean that there will always be a number that is to be compared to another.
The basic formula or equation is always: Percentage = part/whole

Eg 1 If an item is on sale for $15 less than the usual $88, what is the percentage of the discount? Here the part is the $15 and the whole is $88 because you want to see what part $15 is of the whole value $88.
percentage=part/whole
percentage= $15/$88 = .17 = 17/100 = 17 % (This result is rounded to the nearest percentage.)
The “part” can be greater than the “whole”. That will be true whenever the percentage is greater than 100%. The goal for selling cookies was 2500 boxes. 2750 boxes were actually sold. What is the percentage of the goal that was actually sold. Here the part is 2750 because we want to compare it to the goal of 2500. The 2500 is the whole or standard amount. The percentage = 2750/2500
percentage=part/whole
percentage= 2750 boxes/2500 boxes = 1.1 = 110/100 =110%

The word percent means per hundred. Percents are just ratios that allow us to compare things by making the ratio a standard ratio of something/100 or per one hundred.
Which is the biggest discount getting $12 off a $55 purchase or $18 off a 90 purchase? If you rewrite both ratios in a standard form, a percentage which rewrites them as a ratio of x/100, you can compare them.
Percentage 1= 12/55 = .22 = 22/100 = 22%
Percentage 2 = 18/90 = .22 = 20/100 = 20%
By making the ratios more standard by representing them as x/100 or a percentage, we can see that the discount is greater in the first example.

In percentage problems, you are given 2 of the following 3 items:
the percentage
the part
the whole
Write the basic equation, substitute in what you know, and solve for what remains unknown.

OK. For your problem:
You know the percentage and the PART for one calculation. 200 is the part because it tells you that it is 15% of some whole number. The whole is the unknown.
Solve for the unknown. (This one shows all steps.)
pct=part/whole
15%=200/x
x*15%=200/x*x
x*15%=200*x/x
x*15%=200*1
x*15%=200
x*15%*1/15%=200*1/15%
x*1=200/15%
x=200/15%
x=$1333.33
This answers the question, $200 is 15% of what?
The problem asks what is 85% of that amount?

Solve for the unknown. This one doesn't list the steps as above
pct=part/whole
85%=x/1333.33
85%*$1333.33 = x
x = $1133.33


OK. You will learn about percentages from all of the above.
You could also have solved this problem seeing it as ratio problem.
"15% is to 200 as 85% is to x" gives you the ratios:
15%/200=85%/x
15%*x=85%*200
x=(85%*200)/15%
x=$1133.33








Answer by Alan3354(69443)

if 15% is 200.00 then how much is 85%?
----------
200 = Total*0.15
Total = 200/0.15
85% = 0.85*Total = 0.85*200/0.15
85% =~ 1133.33